# Information-Theoretic Bayesian Optimization for Bilevel Optimization Problems

### v1

@@ Abstract
A bilevel optimization problem consists of two optimization problems nested as an upper- and a lower-level problem, in which the optimality of the lower-level problem defines a constraint for the upper-level problem. This paper considers Bayesian optimization (BO) for the case that both the upper- and lower-levels involve expensive black-box functions. Because of its nested structure, bilevel optimization has a complex problem definition and, compared with other standard extensions of BO such as multi-objective or constraint settings, it has not been widely studied. We propose an information-theoretic approach that considers the information gain of both the upper- and lower-optimal solutions and values. This enables us to define a unified criterion that measures the benefit for both level problems, simultaneously. Further, we also show a practical lower bound based approach to evaluating the information gain. We empirically demonstrate the effectiveness of our proposed method through several benchmark datasets.
@@ 1 Introduction
The bilevel optimization is a standard formulation for a decision making problem that has a hierarchical structure. It consists of two optimization problems nested as an upper- and a lower-level problem, in which the optimality of the lower-level problem defines a constraint for the upper-level problem. For example, in the computational materials design, a target property should be optimized under the constraint of the energy minimization. Bilevel optimization techniques is applicable to hierarchical decision makings in a variety of contexts such as inverse optimal control (Suryan et al.,, 2016), chemical reaction optimization (Abbassi et al.,, 2021), and shape optimization (Herskovits et al.,, 2000).
@@
We particularly focus on the case both level problems are defined by expensive black-box functions. Most of BO studies for bilevel optimization consider applying BO only to the upper-level problem (e.g., Kieffer et al.,, 2017; Dogan and Prestwich,, 2023) as pointed out by (Chew et al.,, 2025). Typically, under a selected query for the upper-level problem, repeated queries to the lower-level problem is required, and further, the gradient of the lower-level problem is often assumed (e.g., Fu et al.,, 2024). Islam et al., (2018) and Wang et al., (2021) consider BO in both levels, but repeated queries on lower-level is still required. These approaches are not fully suitable when both levels are expensive black-boxes in which the gradient is not available. On the other hand, recently a few methods without those limitations have also been studied. Ekmekcioglu et al., (2024) combine the Thompson sampling on the upper-level query and a knowledge gradient-based extension of multi-task BO on the lower-level, but the theoretical justification for the combination of these two different criteria has not been revealed. Further, Chew et al., (2025) propose the well-known GP upper confidence bound (UCB) based approach to bilevel BO, called BILBO. Although BILBO has a theoretical regret guarantee, in general, the performance of GP-UCB based methods depend on the selection of the balancing parameter of the exploitation and exploration, because the theoretically recommended value often does not provide the best performance (Srinivas et al.,, 2010).
@@
We propose an information-theoretic approach that considers the simultaneous information gain for both the upper- and lower-optimal solutions and values, which we call bilevel information gain. This enables us to define a unified criterion that measures the benefit for both level problems simultaneously, which is not necessarily common in the case of bilevel methods as mentioned in the previous paragraph. Although the effectiveness of information-theoretic BO has been shown in several different contexts (e.g., Hennig and Schuler,, 2012; Hernández-Lobato et al.,, 2014; Hoffman and Ghahramani,, 2015; Wang and Jegelka,, 2017; Hernández-Lobato et al.,, 2015; Hernandez-Lobato et al.,, 2016; Suzuki et al.,, 2020; Takeno et al., 2022a, ; Takeno et al., 2022b, ; Hvarfner et al.,, 2022; Tu et al.,, 2022), it has not been combined with bilevel optimization, to our knowledge. We first define bilevel information gain by extending the idea of the joint entropy search (Hvarfner et al.,, 2022; Tu et al.,, 2022). Unfortunately, the original definition of bilevel information gain is computationally intractable, and we show that a natural extension of the truncation based approximation, which has been widely employed in information-theoretic BO (e.g., Wang and Jegelka,, 2017), can be derived. By combining the truncation based approximation and a variational lower bound (Takeno et al., 2022b, ), we obtain our criterion called Bilevel optimization via Lower-bound based Joint Entropy Search (BLJES). Further, while we mainly consider ‘coupled setting’ in which upper- and lower-level observations are obtained simultaneously, ‘decoupled setting’, in which a separate observation for each level is available, is also discussed. For example, in the case that the objective function values are outputs of some simulators (e.g., physical simulation), if a common simulator provides both level observations simultaneously, coupled setting is suitable, while if the upper- and the lower-level observations are from different simulators, decoupled setting can be more appropriate. We further propose an extension for the case that each level problem has inequality constraints (i.e., each level problem is a constraint problem).
@@ 2 Preliminaries
$\{\sigma_{t}^{h}(\boldsymbol{x},\boldsymbol{\theta})\}^{2}$ $=k^{h}((\boldsymbol{x},\boldsymbol{\theta}),(\boldsymbol{x},\boldsymbol{\theta}))-\boldsymbol{k}^{h\top}(\boldsymbol{K}^{h}+\{\sigma_{\mathrm{noise}}^{h}\}^{2}\boldsymbol{I})^{-1}\boldsymbol{k}^{h},$
where $\boldsymbol{y}^{h}=\left(y_{1}^{h},\dots,y_{n}^{h}\right)^{\top}$, $\boldsymbol{k}^{h}=\left(k^{h}((\boldsymbol{x},\boldsymbol{\theta}),(\boldsymbol{x}_{1},\boldsymbol{\theta}_{1})),\dots,k^{h}((\boldsymbol{x},\boldsymbol{\theta}),(\boldsymbol{x}_{t},\boldsymbol{\theta}_{n}))\right)^{\top}$, and $\boldsymbol{K}^{h}\in\mathbb{R}^{n\times n}$ is the kernel matrix with an entry $k^{h}((\boldsymbol{x}_{i},\boldsymbol{\theta}_{i}),(\boldsymbol{x}_{j},\boldsymbol{\theta}_{j}))$ at a position $(i,j)$. Here, $\boldsymbol{I}\in\mathbb{R}^{n\times n}$ denotes the identity matrix.
@@ 3.1 Lower Bound of Mutual Information
$\mathrm{MI}(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\ ;\ f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*}\mid{\cal D}_{t}).$
@@
This criterion naturally allows simultaneous consideration of both the upper- and the lower-objectives. Since the direct evaluation of this MI is difficult, we employ a lower bound based approximation. Let $\Omega\coloneqq\{y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})},f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*}\}$. Our lower bound of the MI is derived by a technique that is often used in the context of the variational approximation (e.g., Poole et al.,, 2019):
$\mathrm{MI}(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\ ;\ f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*}\mid{\cal D}_{t})=\mathbb{E}_{\Omega}\left[\log\frac{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},{\cal D}_{t})}{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right]$
$\qquad=\mathbb{E}_{f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*}}\left[\mathbb{E}_{y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},{\cal D}_{t}}\biggl[\log\frac{q(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},{\cal D}_{t})}{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right]$
$\qquad\qquad+\mathrm{KL}\left(p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},{\cal D}_{t})\parallel q(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},{\cal D}_{t})\right)\biggr]$
@@
$\qquad\geq\mathbb{E}_{\Omega}\left[\log\frac{q(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},{\cal D}_{t})}{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right]\eqqcolon\mathrm{LB}(\boldsymbol{x},\boldsymbol{\theta}),$ (3)
@@
where $\mathrm{KL}$ is Kullback-Leibler (KL) divergence and $q(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},{\cal D}_{t})$ is a variational distribution ($q$ can be any density function as far as the KL divergence can be defined). The inequality of the last line can be taken because the KL divergence is non-negative (the equality holds when $p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},{\cal D}_{t})=q(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},{\cal D}_{t})$). Similar lower bounds of the MI have been used in information-theoretic multi-objective and constraint BO (Ishikura and Karasuyama,, 2025; Takeno et al., 2022b, ).
@@
The variational distribution $q$ is an approximation of $p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},{\cal D}_{t})$ for which an exact analytical representation is difficult to know. The difficulty is in the conditioning by the optimal solutions and values, for which the most widely accepted approach in information-theoretic BO is to use truncated distributions (e.g., Wang and Jegelka,, 2017; Suzuki et al.,, 2020; Tu et al.,, 2022). For example, in the case of well-known max-value entropy search (MES), proposed by (Wang and Jegelka,, 2017) for the standard single-level problem $\max_{\boldsymbol{x}}f(\boldsymbol{x})$, the predictive distribution conditioning on the max-value $f_{\rm unc}^{*}\coloneqq\arg\max_{\boldsymbol{x}}f(\boldsymbol{x})$ is approximated by the truncated normal distribution, i.e., $p(f(\boldsymbol{x})\mid f_{\rm unc}^{*})\approx p(f(\boldsymbol{x})\mid f(\boldsymbol{x})\leq f_{\rm unc}^{*})$. When $f_{\rm unc}^{*}$ is given, $f(\boldsymbol{x}^{\prime})\leq f_{\rm unc}^{*}$ should hold for any $\boldsymbol{x}^{\prime}$ (and there should exist at least one $\boldsymbol{x}^{\prime}$ such that $f(\boldsymbol{x}^{\prime})=f_{\rm unc}^{*}$), while MES simplifies this condition so that $f(\boldsymbol{x})\leq f_{\rm unc}^{*}$ holds only for the current $\boldsymbol{x}$. Similar simplifications have been employed by most of information-theoretic BO algorithms and shown superior performance.
@@
$q(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},{\cal D}_{t})$ $\coloneqq p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f{(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))}\leq f^{*},g(\boldsymbol{x}^{*},\boldsymbol{\theta})\leq g^{*},{\cal D}_{t}^{+}),$ (4)
@@
  • When $f^{*}$ is given, $f(\boldsymbol{x}^{\prime},\boldsymbol{\theta}^{*}(\boldsymbol{x}^{\prime}))\leq f^{*}$ should hold for $\forall\boldsymbol{x}^{\prime}$. However, this condition is computationally intractable as mentioned for the case of MES. Based on a similar idea of MES, the condition $f(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))\leq f^{*}$ is only imposed on the current $\boldsymbol{x}$.
@@
  • In the right hand side of (4), ${\cal D}_{t}$ is replaced ${\cal D}_{t}^{+}$. This condition can impose that the GPs satisfy $f(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})=f^{*}$ and $g(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})=g^{*}$ by adding $(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},f^{*},g^{*})$ into the training data.
@@
$\mathrm{LB}(\boldsymbol{x},\boldsymbol{\theta})$ $=\mathbb{E}_{\Omega}\left[\log\frac{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f{(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))}\leq f^{*},g(\boldsymbol{x}^{*},\boldsymbol{\theta})\leq g^{*},{\cal D}_{t}^{+})}{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right]$
$=\mathbb{E}_{\Omega}\left[\log\frac{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))\leq f^{*},{\cal D}_{t}^{+})}{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}+\log\frac{p(y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid g(\boldsymbol{x}^{*},\boldsymbol{\theta})\leq g^{*},{\cal D}_{t}^{+})}{p(y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right].$ (5)

### v2

@@ Abstract
A bilevel optimization problem consists of two optimization problems nested as an upper- and a lower-level problem, in which the optimality of the lower-level problem defines a constraint for the upper-level problem. This paper considers Bayesian optimization (BO) for the case that both the upper- and lower-levels involve expensive black-box functions. Because of its nested structure, bilevel optimization has a complex problem definition, by which bilevel BO has not been widely studied compared with other standard extensions of BO such as multi-objective or constraint problems. We propose an information-theoretic approach that considers the information gain of both the upper- and lower-optimal solutions and values. This enables us to define a unified criterion that measures the benefit for both level problems, simultaneously. Further, we also show a practical lower bound based approach to evaluating the information gain. We empirically demonstrate the effectiveness of our proposed method through several benchmark datasets.
@@ 1 Introduction
The bilevel optimization is a standard formulation for a decision making problem that has a hierarchical structure. It consists of two optimization problems nested as an upper- and a lower-level problem, in which the optimality of the lower-level problem defines a constraint for the upper-level problem. Bilevel optimization techniques is applicable to hierarchical decision makings in a variety of contexts such as inverse optimal control (Suryan et al.,, 2016), chemical reaction optimization (Abbassi et al.,, 2021), and shape optimization (Herskovits et al.,, 2000).
We particularly focus on the case that both level problems are defined by expensive black-box functions, while most of existing studies assume that the lower level problem is not expensive to observe. For example, consider the case that the upper- and lower-objective functions are defined through simulators of a subject of interest, which can occur in a variety of scientific, engineering, and industrial fields. If these simulators consist of expensive computations (such as quantum-mechanical calculations), both level problems are expensive to observe. An example is that the simulator-based optimization of a physical property of inorganic crystals (computational materials design) under the stability constraint (energy minimization) can be seen as an instance of this class of problems.
Most of existing BO studies for bilevel optimization applies BO only to the upper-level problem (e.g., Kieffer et al.,, 2017; Dogan and Prestwich,, 2023) as pointed out by (Chew et al.,, 2025). Typically, for a selected query to the upper-level problem, repeated queries to the lower-level problem is required, and/or further, the gradient of the lower-level problem is often assumed (e.g., Fu et al.,, 2024). A few studies (Islam et al.,, 2018; Wang et al.,, 2021) consider BO in both levels, but repeated queries on the lower-level is still required. These approaches are not fully suitable when both level problems are expensive black-boxes in which the gradient is not available. It is well-known that hyper-parameter optimization (HPO) of a machine learning model can be seen as a bilevel problem. However, the lower objective is usually inexpensive white-box function, and therefore, HPO is different from our focus, while there exist specific studies for HPO bilevel problems (e.g., Lorraine et al.,, 2020). Further, some recent studies assume that the objectives are defined as an average (expectation) of some base functions and the stochastic samples of it are available, which can occur in problems such as HPO, federated learning, and adversarial attacks (e.g., Aghasi and Ghadimi,, 2025; Nazari et al.,, 2025; Qiu et al.,, 2023; Jiao et al.,, 2025), while our framework do not employ this assumption (further, in these papers, objectives are not necessarily expensive).
@@
On the other hand, recently a few methods have been studied for the both level expensive setting. Ekmekcioglu et al., (2024) combine the Thompson sampling on the upper-level query and a knowledge gradient-based extension of multi-task BO on the lower-level, but a rationale behind the combination of two different criteria has not been revealed. Further, Chew et al., (2025) propose the well-known GP upper confidence bound (UCB) based approach to bilevel BO, called BILBO. Although BILBO has a theoretical regret guarantee, in general, the performance of GP-UCB based methods depend on the selection of the balancing parameter of the exploitation and exploration, because the theoretically recommended value often does not provide the best performance (Srinivas et al.,, 2010).
@@
We propose an information-theoretic approach that considers the simultaneous information gain for both the upper- and lower-optimal solutions and values, which we call bilevel information gain. This enables us to define a unified criterion that measures the benefit for both level problems simultaneously, which is not necessarily common in the case of bilevel methods as already mentioned. Although the effectiveness of information-theoretic BO has been shown in several different contexts (e.g., Hennig and Schuler,, 2012; Hernández-Lobato et al.,, 2014; Hoffman and Ghahramani,, 2015; Wang and Jegelka,, 2017; Hernández-Lobato et al.,, 2015; Hernandez-Lobato et al.,, 2016; Suzuki et al.,, 2020; Takeno et al., 2022a, ; Takeno et al., 2022b, ; Hvarfner et al.,, 2022; Tu et al.,, 2022), it has not been combined with bilevel optimization, to our knowledge. We first define bilevel information gain by extending the idea of the joint entropy search (Hvarfner et al.,, 2022; Tu et al.,, 2022). Unfortunately, the original definition of bilevel information gain is computationally intractable, and we show that a natural extension of the truncation based approximation, which has been widely employed in information-theoretic BO (e.g., Wang and Jegelka,, 2017), can be derived. By combining the truncation based approximation and a variational lower bound (Takeno et al., 2022b, ), we obtain our criterion called Bilevel optimization via Lower-bound based Joint Entropy Search (BLJES).
Further, while we mainly consider ‘coupled setting’ in which upper- and lower-level observations are obtained simultaneously, ‘decoupled setting’, in which a separate observation for each level is available, is also discussed. For example, in the case that the objective function values are outputs of some simulators (e.g., physical simulation), if a common simulator provides both level observations simultaneously, coupled setting is suitable, while if the upper- and the lower-level observations are from different simulators, decoupled setting can be more appropriate. We further propose an extension for the case that each level problem has inequality constraints (i.e., each level problem is a constraint problem).
@@ 2 Preliminaries
$\{\sigma_{t}^{h}(\boldsymbol{x},\boldsymbol{\theta})\}^{2}$ $=k^{h}((\boldsymbol{x},\boldsymbol{\theta}),(\boldsymbol{x},\boldsymbol{\theta}))$
$-\boldsymbol{k}^{h\top}(\boldsymbol{K}^{h}+\{\sigma_{\mathrm{noise}}^{h}\}^{2}\boldsymbol{I})^{-1}\boldsymbol{k}^{h},$
where $\boldsymbol{k}^{h}\!=\!(k^{h}((\boldsymbol{x},\boldsymbol{\theta}),(\boldsymbol{x}_{1},\boldsymbol{\theta}_{1})),\dots,k^{h}((\boldsymbol{x},\boldsymbol{\theta}),(\boldsymbol{x}_{t},\boldsymbol{\theta}_{n})))^{\top}$, $\boldsymbol{y}^{h}=\left(y_{1}^{h},\dots,y_{n}^{h}\right)^{\top}$, and $\boldsymbol{K}^{h}\in\mathbb{R}^{n\times n}$ is the kernel matrix with an entry $k^{h}((\boldsymbol{x}_{i},\boldsymbol{\theta}_{i}),(\boldsymbol{x}_{j},\boldsymbol{\theta}_{j}))$ at a position $(i,j)$. Here, $\boldsymbol{I}\in\mathbb{R}^{n\times n}$ denotes the identity matrix.
@@ 3.1 Lower Bound of Mutual Information
$\mathrm{MI}(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\ ;\ o^{*}\mid{\cal D}_{t}),$
@@
for which an illustration is shown in Fig. 2. This criterion naturally allows simultaneous consideration of both the upper- and the lower-objectives. Since the direct evaluation of this MI is difficult, we employ a lower bound based approximation. Let $\Omega\coloneqq\{y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})},f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*}\}$. Our lower bound of the MI is derived by a technique that is often used in the context of the variational approximation (e.g., Poole et al.,, 2019):
$\mathrm{MI}(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\ ;\ o^{*}\mid{\cal D}_{t})$
$=\mathbb{E}_{\Omega}\left[\log\frac{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid o^{*},{\cal D}_{t})}{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right]$
$=\mathbb{E}_{o^{*}}\left[\mathbb{E}_{y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid o^{*},{\cal D}_{t}}\biggl[\log\frac{q(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid o^{*},{\cal D}_{t})}{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right]$
$\ \ \ \ +\mathrm{KL}\left(p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid o^{*},{\cal D}_{t})\parallel q(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid o^{*},{\cal D}_{t})\right)\biggr]$
@@
$\geq\mathbb{E}_{\Omega}\left[\log\frac{q(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid o^{*},{\cal D}_{t})}{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right]\eqqcolon\mathrm{LB}(\boldsymbol{x},\boldsymbol{\theta}),$ (3)
@@
where $\mathrm{KL}$ is Kullback-Leibler (KL) divergence and $q(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid o^{*},{\cal D}_{t})$ is a variational distribution ($q$ can be any density function as far as the KL divergence can be defined). The inequality of the last line can be taken because the KL divergence is non-negative (the equality holds when $p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid o^{*},{\cal D}_{t})=q(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid o^{*},{\cal D}_{t})$). Similar lower bounds of the MI have been used in information-theoretic multi-objective and constraint BO (Ishikura and Karasuyama,, 2025; Takeno et al., 2022b, ).
Figure 2: Schematic illustration of random variables in $\mathrm{MI}(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\ ;\ o^{*}\mid{\cal D}_{t})$. The left panel represents input space of $f$ and $g$ (heatmap is the true functions), and the right panel is the output space. The predictions at current $(\boldsymbol{x},\boldsymbol{\theta})$ are $y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}$ and $y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}$, depicted as the green distribution in the output space. At the optimal solution $(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$, which is also a random variable defined through the GP posteriors, the objective function values are $f^{*}=f(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$ and $g^{*}=g(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$. The optimal values $f^{*}$ and $g^{*}$ are represented as the blue distribution in the output space. The optimal solution $(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$, represented as the red distribution in the input space, should exist on $\boldsymbol{\theta}^{*}(\boldsymbol{x})$, i.e., the optimal point of the lower-level problem. The red and blue stars are ‘samples’ of the optimal solutions $(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$ and values $(f^{*},g^{*})$, respectively.
@@
The variational distribution $q$ is an approximation of $p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid o^{*},{\cal D}_{t})$ for which an exact analytical representation is difficult to know. The difficulty is in the conditioning by the optimal solutions and values, for which the most widely accepted approach in information-theoretic BO is to use truncated distributions (e.g., Wang and Jegelka,, 2017; Suzuki et al.,, 2020; Tu et al.,, 2022). For example, in the case of well-known max-value entropy search (MES), proposed by (Wang and Jegelka,, 2017) for the standard single-level problem $\max_{\boldsymbol{x}}f(\boldsymbol{x})$, the predictive distribution conditioning on the max-value $f_{\rm unc}^{*}\coloneqq\arg\max_{\boldsymbol{x}}f(\boldsymbol{x})$ is approximated by the truncated normal distribution, i.e., $p(f(\boldsymbol{x})\mid f_{\rm unc}^{*})\approx p(f(\boldsymbol{x})\mid f(\boldsymbol{x})\leq f_{\rm unc}^{*})$. When $f_{\rm unc}^{*}$ is given, $f(\boldsymbol{x}^{\prime})\leq f_{\rm unc}^{*}$ should hold for any $\boldsymbol{x}^{\prime}$ (and there should exist at least one $\boldsymbol{x}^{\prime}$ such that $f(\boldsymbol{x}^{\prime})=f_{\rm unc}^{*}$), while MES simplifies this condition so that $f(\boldsymbol{x})\leq f_{\rm unc}^{*}$ holds only for the current $\boldsymbol{x}$. Similar simplifications have been employed by most of information-theoretic BO algorithms and shown superior performance.
@@
$q(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid o^{*},{\cal D}_{t})$
$\coloneqq p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f{(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))}\leq f^{*},g(\boldsymbol{x}^{*},\boldsymbol{\theta})\leq g^{*},{\cal D}_{t}^{+}),$ (4)
@@
  • When $f^{*}$ is given, $f(\boldsymbol{x}^{\prime},\boldsymbol{\theta}^{*}(\boldsymbol{x}^{\prime}))\leq f^{*}$ should hold for $\forall\boldsymbol{x}^{\prime}$. However, this condition is computationally intractable as mentioned for the case of MES. Based on a similar idea to MES, the condition $f(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))\leq f^{*}$ is only imposed on the current $\boldsymbol{x}$.
@@
  • In the right hand side of (4), ${\cal D}_{t}$ is replaced with ${\cal D}_{t}^{+}$. This condition imposes that the GPs satisfy $f(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})=f^{*}$ and $g(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})=g^{*}$ by adding a noiseless observation $(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*},f^{*},g^{*})$ into the training data.
@@
$\mathrm{LB}(\boldsymbol{x},\boldsymbol{\theta})=$
${\small\mathbb{E}_{\Omega}\!\left[\log\frac{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f{(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))}\leq f^{*},g(\boldsymbol{x}^{*},\boldsymbol{\theta})\leq g^{*},{\cal D}_{t}^{+})}{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right]}$
$=\mathbb{E}_{\Omega}\Biggl[\log\frac{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))\leq f^{*},{\cal D}_{t}^{+})}{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}$
$\qquad\qquad+\log\frac{p(y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid g(\boldsymbol{x}^{*},\boldsymbol{\theta})\leq g^{*},{\cal D}_{t}^{+})}{p(y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\Biggr].$ (5)